### v2

# Improving Retrieval Augmented Language Model with Self-Reasoning

Abstract
The Retrieval-Augmented Language Model (RALM) has shown remarkable performance on knowledge-intensive tasks by incorporating external knowledge during inference, which mitigates the factual hallucinations inherited in large language models (LLMs). Despite these advancements, challenges persist in the implementation of RALMs, particularly concerning their reliability and traceability. To be specific, the irrelevant document retrieval may result in unhelpful response generation or even deteriorate the performance of LLMs, while the lack of proper citations in generated outputs complicates efforts to verify the trustworthiness of the models. To this end, we propose a novel self-reasoning framework aimed at improving the reliability and traceability of RALMs, whose core idea is to leverage reasoning trajectories generated by the LLM itself. The framework involves constructing self-reason trajectories with three processes: a relevance-aware process, an evidence-aware selective process, and a trajectory analysis process. We have evaluated our framework across four public datasets (two short-form QA datasets, one long-form QA dataset, and one fact verification dataset) to demonstrate the superiority of our method, which can outperform existing state-of-art models and can achieve comparable performance with GPT-4, while only using 2,000 training samples.
Improving Retrieval Augmented Language Model with Self-Reasoning
Yuan Xia, Jingbo Zhou, Zhenhui Shi, Jun Chen, Haifeng Huang Baidu Inc., China. {xiayuan,zhoujingbo,shizhenhui,chenjun22,huanghaifeng}@baidu.com
1 Introduction
Figure 1: An example of how self-reasoning framework generates reasoning trajectories with a relevant-aware process, an evidence-aware selective process, and a trajectory analysis process.
The Retrieval-Augmented Language Model (RALM) has become a critical enhancement for Large Language Models (LLMs) by integrating external knowledge during inference. Despite their advanced capabilities in language understanding and generation Brown et al. (2020); Touvron et al. (2023), LLMs are prone to producing hallucinated and inaccurate content, particularly in knowledge-intensive tasks Ji et al. (2023). Augmenting LLMs with relevant information obtained from external sources like Wikipedia and search engines has proven effective in reducing these inaccuracies Guu et al. (2020); Lewis et al. (2020); Borgeaud et al. (2022); Izacard et al. (2022); Asai et al. (2024). This method has proven effective in mitigating the factual hallucinations inherent in LLMs Kwiatkowski et al. (2019); Petroni et al. (2021); Ram et al. (2023).
Nevertheless, there are still limitations associated with RALMs, particularly in terms of reliability and traceability. First, the reliability of the retrieved information remains a significant concern. Previous studies have shown that noisy retrieval can adversely affect the performance of an LLM Menick et al. (2022); Li et al. (2023), as irrelevant data can lead to misguided responses and disturb the model’s ability to leverage its intrinsic knowledge effectively. Second, the interpretability and traceability of outputs generated by RALMs need to be improved. While RALMs incorporate retrieved documents during both training and inference phases, they may fail to cite these documents explicitly,
thus complicating the process of tracing and verifying the claims made by LLMs. To improve the retrieval robustness, recent studies have explored incorporating natural language inference (NLI) models Honovich et al. (2022) and document summarization models during inference Yoran et al. (2023); Xu et al. (2023a). However, the effectiveness of these external NLI and summarization models largely affects the overall performance of RALMs. The training and optimization of these auxiliary models require additional costs. Consequently, determining the most appropriate training and selection methods for these NLI and summarization models presents a further challenge in the application of such approaches.
To address the above limitations, we propose a novel end-to-end self-reasoning framework to improve the performance of RALMs. Our intuition is that the explicit self-reasoning trajectory crafted by LLMs can improve the retrieval robustness and accuracy of question answering. During the pre-training phase, an LLM primarily focuses on knowledge acquisition, yet it does not learn to reason from retrieved documents to generate answers. To address this, a feasible approach is to incorporate reasoning trajectories into a post-training phase. Such an approach could potentially teach the model to reason and distinguish relevant and irrelevant documents, improving its response accuracy to queries. An example of how our self-reasoning framework generates reasoning trajectories is illustrated in Figure 1. In contrast, as shown in the middle part of Figure 2, the conventional RALM methods gather all documents in a non-selective manner, leading to the distraction of the LLM by irrelevant content and consequently resulting in the generation of erroneous answers.
Our framework comprises three self-reasoning processes: 1) a Relevance-Aware Process (RAP), which instructs the LLM to judge the relevance between the retrieved documents and the question, 2) an Evidence-Aware Selective Process (EAP), which directs the LLM to choose and cite relevant documents, and then automatically select snippets of key sentences as evidence from the cited documents, 3) a Trajectory Analysis Process (TAP), which requires the LLM to generate a concise analysis based on all gathered self-reasoning trajectories generated by previous two processes and subsequently provide the final inferred answer. Furthermore, we propose a gradual training method by employing stage-wise masking strategies to enhance the performance of our framework.
We summarize our contributions as follows:
  • We propose a novel end-to-end self-reasoning framework that improves the robustness of RALMs by leveraging reasoning trajectories generated by the LLM itself, without the need for external models or tools.
  • We carefully design three processes to enhance the interpretability and traceability of RALMs by requiring LLMs to explicitly generate snippets and citations from documents, and further explain the reason why cited documents can help answer the question.
  • We evaluate our framework on four public datasets (two short-form QA, one long-form QA, and one fact verification), demonstrating that our method surpasses existing state-of-art models in performance, achieving this with only using 2,000 training samples.
Figure 2: An illustration of the self-reasoning framework for improving the RALMs. The upper is the basic LLMs which answer the question by inherent knowledge. The middle is the standard retrieval augmented LMs, which use retrieved documents to help answer the question. The bottom is our self-reasoning framework which uses self-generated reason trajectories to output answers.
2 Related Work
2.1 Retrieval-augmented LMs
Many studies have investigated augmenting the performance of LLMs with externally retrieved information Izacard et al. (2022); Guu et al. (2020); Borgeaud et al. (2022) and some of them pre-train language models with retrieved passages. For works focusing on RALMs with citations, Menick et al. (2022); Nakano et al. (2021) instruct or train an LLM to answer questions with retrieved documents while providing citations. Gao et al. (2023b) proposes an end-to-end system to retrieve supporting evidence and generate answers with citations, while only focusing on prompting without updating their model weights. Other works instruct or fine-tune LLMs to use external tools to retrieve dynamically Schick et al. (2023); Yao et al. (2023); Jiang et al. (2023), which offers an adaptive method of when and what to search. Gao et al. (2023a) improves the attribution and factuality of language models by taking outputs of LLMs and applying a post-process retrieve-and-edit approach.
2.2 Robustness for RALMs
To improve the robustness of RALMs, Yoran et al. (2023) use a natural language inference model to filter out irrelevant documents before RALMs, and Xu et al. (2023a) use a memorization model to filter out or compress retrieved documents before using them to prompt an LLM. The concurrent work Yu et al. (2023) generates summaries of documents as notes before RALMs, and Baek et al. (2023) uses a separate small language model as a verifier to detect and correct errors in LLMs during retrieval for generating factually correct outputs. Different from the above works, ours identifies key sentences and cites relevant documents through an end-to-end framework that eliminates irrelevant ones, without dependence on external inference models.
The most similar work to ours is presented by Asai et al. (2024), who developed a method that teaches models to retrieve information using designed reflection tokens. However, this approach needs to train extra critic models and generator models to predict the reflection tokens, which requires tens of thousands of extra training samples. In contrast, our method does not rely on special tokens. Instead, we construct reasoning trajectories and then use them to directly enhance the performance of LLMs, offering a more efficient and scalable solution. More related works on LLMs for reasoning are discussed in the Appendix §$\S$§A.1.
3 Preliminary
We formally define the problem of retrieval augmented generation with self-reasoning. Given a query $q$ and a corpus of documents $\mathcal{D}$, an LLM-generated answer with $m$ statements and $n$ tokens can be defined as $y=(s_{1},s_{2},\cdots,s_{m})=(w_{1},w_{2},\cdots,w_{n})$, where $s_{i}$ is the $i$-th statement and $w_{j}$ is the $j$-th token in the generated answer. In addition, for long-form QA settings, each statement $s_{i}$ should cite a list of documents $C_{i}=\{c^{(1)}_{i},c^{(2)}_{i},...\}$, where $c^{(k)}_{i}\in\mathcal{D}$. In our work, we train an LLM (e.g. LLaMA2) to first generate reasoning trajectories $\tau$ through self-reasoning and then to generate answers ${y}^{*}$ (including long-form answers and short-form answers) on condition of $\tau$. The model output is $y={\rm{concat}}(\tau,y^{*})$, which is the concatenation of $\tau$ and $y^{*}$. Note that the generations of $\tau$ and $y^{*}$ are done in a single pass within our self-reasoning framework.
4 Method
In this section, we provide a detailed implementation of the self-reasoning process and the method to construct reasoning trajectories.
In this work, we propose a novel framework which involves three processes: 1) a Relevance-Aware Process (Sec §$\S$4.1), 2) an Evidence-Aware Selective Process (Sec §$\S$4.2), and 3) a Trajectory Analysis Process (Sec §$\S$4.3). An illustration of our self-reasoning framework for improving the RALMs is shown in Figure 2. In Sec 4.4, we demonstrate the process of data generation and quality control, and Sec 4.5 presents the details of model training.
4.1 Relevance-Aware Process
In this work, we choose the DPR Karpukhin et al. (2020) and the Contriever Izacard et al. (2021) as the default retriever $R$ to recall the top-$k$ relevant documents. When presented with a question and a set of documents, people can determine whether the question is relevant to the retrieved documents. Therefore, we first instruct the model to judge the relevance between the retrieved documents $\mathcal{D}$ and the given question $q$. We further request the model to explicitly generate reasons explaining why given documents are identified as relevant. The output should include two fields as relevant and relevant reason, as depicted in Figure 2. Noting that if all of the retrieved documents are irrelevant, the model should provide an answer based on the internal knowledge acquired during its pre-training phase. We define the self-reasoning trajectories generated by RAP as $\tau_{r}$.
4.2 Evidence-Aware Selective Process
When required to answer the question, people generally will first identify the crucial sentences from the provided documents, and then cite or highlight them as key points. This process of citing the document facilitates reading comprehension and can serve as a technique for combining multiple short answers to address various aspects. While people may carry out this selective process and citation instantaneously, LLMs need to formulate the self-reasoning trajectories explicitly.
In our work, we require the LLM to explicitly state the reason why the selected sentence is supportive and plausible in answering the question. We define the selected sentence as evidence in our paper. Specifically, after retrieving the top-$k$ documents, the self-reasoning method for Evidence-Aware Selective Process can be formulated as follows: First, we instruct the LLM to choose relevant documents and automatically select snippets of key sentences for the selected documents. Then, we request the LLM to output the reason why the selected snippets can answer the question. The intermediate output is a list containing multiple contents, each content should include two fields, as cite content and reason for cite, which is illustrated in Figure 2. We define the self-reasoning trajectories generated by EAP as $\tau_{e}$.
4.3 Trajectory Analysis Process
Finally, we consolidate all the self-reasoning trajectories ($\tau_{r}$ and $\tau_{e}$) in the previous processes together to form a chain of reasoning snippets, thereby enhancing the overall performance of the retrieval augmentation generation. Specifically, we ask the LLM to analyze the reasoning trajectories within itself and ultimately to output a concise analysis and a short answer. We instruct the LLM to output content with two fields as analysis and answer, which is shown in Figure 2. We define the self-reasoning trajectories generated by TAP as $\tau_{a}$. In this work, the analysis output is defined as a long-form answer, and the answer output is defined as a short-short answer. In Section 5.2, we further explored the performance of long-form and short-form QA settings.
4.4 Data Generation and Quality Control
Training Data Generation.
For the Relevance-Aware Process data generation, as manually labeling the relevant and irrelevant documents is label-intensive, we request the GPT-4 OpenAI (2023) to generate answers as ground truth. Specifically, we instruct the GPT-4 to generate labels regarding irrelevant fields, and further to output the reasons why the given documents cannot answer the question. We concatenate the given question and the retrieved documents as positive samples. For negative samples, we randomly select a different question from the training set and retrieve the top-$k$ documents related to it. These documents are then concatenated with the initial question to form negative samples. To avoid order bias in the training data, we shuffle the order of the documents.
For the EAP and TAP data generation, manually annotating the citation of the document and writing the self-reasoning process for each question is not feasible in practice. Therefore, we follow a similar process to RAP, we first instruct the GPT-4 to generate a snippet of selected documents and subsequently output the reasoning process as trajectories. The method to construct the EAP training data is the same as RAP except that the instructions for the GPT-4 are different. The details of the instructions are shown in Appendix §$\S$§A.2.
Data Quality Control.
For training data generation, correct and comprehensive reasoning trajectories are very important. When training an LLM, the quality of the training samples is more important than the quantity Zhou et al. (2023). As we cannot guarantee the correctness of self-reasoning trajectories and citations by the GPT-4, we develop two efficient methods to control the quality of data generation: 1) The first method is to use the off-the-shelf tools  [^1] Tools are available at https://github.com/princeton-nlp/ALCE/tree/main in Gao et al. (2023b) to automatically verify the performance of data generation for document citations. We calculate the citation precision and recall score for each training sample and filter out scores lower than our pre-defined thresholds $\delta_{p}$ and $\delta_{r}$, for citation precision and recall, respectively. 2) Second, though the validation of self-reasoning trajectories and citations generated by GPT-4 is challenging, verifying the correctness of the final answer is straightforward. Therefore, we filter out the trajectories that lead to the incorrect answers and only keep the correct ones. We totally generate 10,000 training samples by GPT-4, after the filtering strategy by quality control, we finally keep 2,000 training samples with high quality.
4.5 Model Training
We train our self-reasoning generation model $\phi$ by our constructed corpus which is augmented with self-reasoning trajectories $\tau$ using the standard language modeling objective, maximizing likelihood:
$\max_{\mathcal{\phi}}\mathbb{E}_{(q,\tau,y)\sim\mathcal{D}_{sr}}\log p_{\mathcal{\phi}}(y\mid\tau,q)p_{\mathcal{\phi}}({\tau\mid q})$ (1)
where $\tau=\tau_{r}\oplus\tau_{e}\oplus\tau_{a}$ are the self-reasoning trajectories, $\oplus$ is the concatenation operator, $\tau_{r},\tau_{e},\tau_{a}$ are trajectories generated by above three processes respectively. $q$ is the provided question, and $y$ is the model output, including the intermediate reason trajectories and the final answer. $\mathcal{D}_{sr}$ is the training corpus augmented with self-reasoning trajectories.
During training, we observed that it is more challenging to ensure the correctness of an LLM with 13B parameters when generating long reasoning trajectories than short ones. We hypothesize that an LLM’s effective reasoning length is limited and exceeding this limit might lead to error accumulation during the inference stage. Therefore, we propose a gradual training method by employing stage-wise masking strategies to gradually learn to generate long trajectories.
Specifically, we propose a stage-wise training process while we train the LLM stage by stage. In the first stage, we mask the trajectories produced by the next two stages (EAP and TAP) and train the model with a learning rate $r_{a}$. Then in the second stage, we only mask the trajectories generated by TAP and train the model with a learning rate $r_{b}$. Finally, we concatenate the reasoning trajectories from all stages and put them into a self-reasoning LLM for end-to-end training with a learning rate $r_{c}$. Hyper-parameters for training are described in Appendix §$\S$§A.4.
5 Experiments
Models NaturalQuestion PopQA FEVER ASQA
(acc) (acc) (acc) (em-recall) (precision) (recall)
Baselines without retrieval
LLaMA27B 19.2 18.4 23.2 10.2 - -
LLaMA213B 24.0 22.6 25.3 15.3 - -
LLaMA27B-chat 20.2 21.5 26.5 16.3 - -
LLaMA213B-chat 23.2 25.9 28.4 18.3 - -
Baselines with retrieval
LLaMA27B 27.8 47.8 39.8 28.5 13.6 9.59
LLaMA213B 34.0 48.1 35.2 26.8 21.8 16.3
LLaMA27B-chat 27.4 52.9 43.4 25.3 34.5 33.2
LLaMA213B-chat 32.7 53.5 53.4 26.4 39.4 38.4
Vicuna7B Chiang et al. (2023) 28.0 55.2 62.4 24.3 45.7 40.8
Vicuna13B Chiang et al. (2023) 35.4 56.1 60.6 27.3 51.3 50.2
LLaMA2-FT7B 36.8 54.4 67.5 28.5 47.2 45.4
ReAct Yao et al. (2023) - - 64.6 - - -
RECOMP Xu et al. (2023a) 38.4 - - - - -
Self-RAG7B Asai et al. (2024) 37.2 54.9 70.2 30.0 66.9 67.8
Self-RAG13B Asai et al. (2024) 38.8 55.8 72.1 31.7 70.3 71.3
Self-Reasoning7B 38.0 54.2 78.6 33.9 66.3 70.8
Self-Reasoning13B 41.4 57.3 83.9 35.2 71.2 72.3
GPT-4 46.6 62.5 87.7 41.3 75.6 68.5
Table 1: Performance comparisons with different baseline models on two short-form QA datasets, a long-form QA dataset, and a fact verification dataset. The numbers with bold black represent the best results excluding GPT-4. The results are averaged over five runs, and presented with standard variance values omitted (all $\leq 2\%$).
5.1 Datasets and Settings
To demonstrate the effectiveness of our proposed self-reasoning framework, we conduct an extensive experimental evaluation on two short-form QA datasets (NaturalQuestion Kwiatkowski et al. (2019) and PopQA Mallen et al. (2023)), one long-form QA dataset (ASQA Stelmakh et al. (2022)), and one fact verification dataset (FEVER Thorne et al. (2018)). Detailed descriptions of the datasets can be found in Appendix §$\S$§A.3. We explore off-the-shelf retrievers. We use the DPR Karpukhin et al. (2020) and the Contriever-MS MARCO Izacard et al. (2021) to retrieve the top five documents from Wikipedia.
By default, we use the DPR as a retriever for the NQ, as the DPR has been fine-tuned on the high-quality NQ data. On the PopQA, where question and answer pairs are created based on Wikipedia in 2022, therefore, for the PopQA, we use the December 2020 preprocessed Wikipedia corpus provided by Izacard et al. (2022) and use the Contriever as a retriever. For the ASQA dataset, we use GTR Ni et al. (2022) as a retrieval that corresponds to the experimental settings in Gao et al. (2023b). More settigns can be found in Appendix §$\S$§A.4.
5.2 Evaluation Metrics
We use different evaluation metrics for short-form QA, long-form QA, and fact verification tasks.
Short-form QA metrics.
We report accuracy for short-form QA tasks, which is based on whether ground-truth answers are included in the model predictions instead of strictly requiring exact matching, following Mallen et al. (2023); Schick et al. (2023).
Long-form QA metrics.
For long-form QA tasks, we report the EM recall as a correctness metric, and the citation recall and the citation precision for citation quality, which are the same as the metrics in Gao et al. (2023b).
Fact verification metrics.
For the fact verification task, we report the accuracy as a metric, which is a three-class classification accuracy, following Thorne et al. (2018).
5.3 Baseline Models
Baseline models without retrieval.
We evaluate strong open-source pre-trained LLMs as baseline models. For basic LLMs, we test LLaMA2-7B, LLaMA2-13B Touvron et al. (2023) and its instruction-tuned chat version LLaMA2-Chat-7B, LLaMA2-Chat-13B.
Baseline models with retrieval.
To assess the performance of the retrieval augmented LMs, we use retrievers described in Section 5.1. First, we benchmark the models using the LLaMA2 and the Vicuna Chiang et al. (2023) series models for baselines. Additionally, for a fair comparison, we also include LLaMA2-FT, where LLaMA2 is fine-tuned on all the training samples generated by GPT-4 except the self-reasoning trajectories. To establish strong baselines, we compare our method against RECOMP Xu et al. (2023a), ReAct Yao et al. (2023), and Self-RAG Asai et al. (2024), all of which are trained with extra GPT-4 generated samples or external tools. We also compare our framework with GPT-4 OpenAI (2023). We include categorical comparisons with the baseline models in the Appendix §$\S$§A.5.
Figure 3: Noise robustness experiment results on three different datasets: (a) On the left is the NQ dataset, (b) in the middle is the PopQA dataset, (c) on the right is the FEVER dataset. The Self-RAG and Vicuna are 13B parameter size models.
5.4 Main Results
Table 1 shows the performance comparisons with different methods on the four public datasets. For short-form QA evaluations, the performance of LLMs with augmented retrieval is consistently better than that of basic ones, affirming the effectiveness of the augmented approach. Notably, under the same order of magnitude parameters, our self-reasoning framework outperforms most of the strong baseline LLMs. Specifically, compared to the Self-RAG, our framework is an end-to-end system trained with only 2,000 self-reasoning trajectory samples. In contrast, the Self-RAG requires training additional critic LMs to predict reflection tokens using an additional 46,000 instances generated by GPT-4. This efficiency not only simplifies the training process but also significantly reduces resource consumption.
In the context of long-form QA evaluations, for the metrics of EM recall, it needs to comprehend multiple documents and merge answers. The EAP and TAP are specifically designed for multi-document reading comprehension, enabling our performance to surpass other baselines. In terms of citation evaluation metrics, our self-reasoning framework can achieve better results than GPT-4 in ASQA citation recall metrics (72.3 vs. 68.5).
This is largely due to the reasoning trajectories generated in the EAP, which can enhance the recall and precision of citation evaluation, leading to more interpretable generations.
For fact verification evaluations, we observed that self-seasoning is dominantly superior to all baseline models. Our method achieves a much higher accuracy rate than the Self-RAG model (83.9 vs. 72.1). The RAP in our framework is designed to judge the relevance between the retrieved documents and the question, which leads to the notable enhancement in accuracy for this fact verification task.
To clearly demonstrate the practical applications and benefits of our self-reasoning framework, we provide a case study for a more in-depth analysis in the Appendix §$\S$§A.6, which illustrates how the framework operates in real-world scenarios.
Models NQ PopQA FEVER
(acc) (acc) (acc)
Origin 41.4 57.3 83.9
[…]
w/o (EAP) 37.2 53.2 78.4
w/o (TAP) 38.2 53.4 81.2
[…]
w/o (QC) 37.7 54.2 80.8
Table 2: The ablation study on two short-form QA datasets and a fact verification dataset with 13B parameter size models. In the table, the origin represents our self-reasoning model enhanced with self-generated trajectories.
6 Analysis
6.1 Ablation Study
We conduct an ablation study on two short-form QA datasets and a fact verification dataset to analyze the individual contributions of each process within our proposed self-reasoning framework. We further explore the effectiveness of the gradual learning (GL) method and the quality control (QC) of data generation (a detailed analysis described in Appendix §$\S$§A.7). The main ablation study results are shown in Table 2.
Effectiveness of Relevant-Aware Process
First, we evaluate the effect of the Relevant-Aware Process (RAP). The removal of the RAP causes the overall performance to drop in two short-form QA datasets and a fact verification dataset, suggesting that preliminary consideration of the relevance between questions and retrieved documents can help improve performance. We notice that the performance declines most significantly in the FEVER dataset. Detecting irrelevant documents is critical in the fact-verification task. Our model will immediately output NotEnoughInfo if it detects that all documents are irrelevant.
Effectiveness of Evidence-Aware Process
Then we evaluate the effect of the Evidence-Aware Selective Process (EAP). Removing the EAP causes the overall performance of the average accuracy to decline from 60.9 to 56.3 in three short-form QA datasets. This reduction indicates that snippets of key sentences and document citations generated through self-reasoning are instrumental in boosting accuracy.
Effectiveness of Trajectory Analysis Process
Finally, we evaluate the effect of the Trajectory Analysis Process (TAP). When excluding the TAP, we can observe a performance decline on all three datasets, demonstrating that self-analysis based on two previous processes generated trajectories can also improve the performance of LLMs. Note that the analysis content generated by TAP is indispensable for the long-form QA evaluation.
6.2 Retrieval Robustness Analysis
As retrievers are not perfect and past work has shown that noisy retrieval can have negative effects on the performance of LLMs Petroni et al. (2020); Li et al. (2023). In this section, we design two kinds of settings to validate the robustness of RALMs. In the first setting, we test whether the order of the retrieved documents will affect the performance of the RALMs. Specifically, after retrieving the top-$k$ documents using retrievals (such as the DPR) with a descending relevance score, we randomly shuffle the order of the retrieved documents and then input them to an LLM. In the second setting, we test how noisy documents impact the performance of LLMs. When retrieving the top-$k$ documents from the given question, we randomly replace 50% of the retrieved documents with other documents sampled from a different question in the dataset.
Figure 3 shows the noise robustness experiment results on three datasets. Our self-reasoning framework consistently outperforms the Self-RAG and Vicuna models. We observe that random shuffling of retrieved documents has a minimal impact on the performance of RALMs. If the provided documents are supportive, it is trivial for a RALM to determine the correct answer. However, when presented with noisy documents, all models experience a decline in performance. The performance drop for our self-reasoning framework is relatively minimal, which demonstrates that our proposed method is robust even when dealing with noisy documents.
Figure 4: Human citation quality evaluation vs. automatic citation evaluation on the long-form ASQA dataset.
6.3 Citation analysis
As the automatic evaluation by the NLI model cannot detect partially supported citations, we discuss the analysis of citations with human evaluation in this section. Similarly to Liu et al. (2023), we conduct a human evaluation on two dimensions: 1) citation recall: annotators are given a statement and all documents that the statement refers to and are asked to judge whether the documents fully support the given statement; 2) citation precision: given a statement and one of its citations, annotators are asked to validate whether the citation fully supports, partially supports or does not support the statement. Each citation gets a precision score of 1 if the output sentence has a citation recall of 1 and this citation at least partially supports it. Details of human annotation can be found in Appendix §$\S$§A.8.
As shown in Figure 4, we observe that the relative rankings by human evaluation align well with those from the automatic evaluation, and the human evaluation often yields a closely higher score when compared with the automatic evaluation.
7 Conclusion
RALMs can effectively enhance the performance of LLMs in handling knowledge-intensive tasks. Despite their effectiveness, notable concerns about their reliability and traceability persist. To address these limitations, we propose a novel self-reasoning framework to improve the performance of RALMs by using reasoning trajectories generated by the LLM itself. It is comprised of a relevance-aware process, an evidence-aware selective process, and a trajectory analysis process. We conduct extensive experiments on four public datasets to demonstrate the superiority of our framework over existing state-of-the-art models.
8 Limitations
In this work, we mainly focus on improving the performance of RALMs with a self-reasoning framework on the task of open domain question answering and fact verification. Although we believe our framework can involve the distribution of real-world user questions, as we have evaluated in four public datasets, we acknowledge that we have not explored more challenging scenarios, such as multi-hop reasoning, code generation, and arithmetic reasoning. In future work, more challenging reasoning tasks, such as arithmetic reasoning, should be explored for the self-reasoning framework. We believe that our framework can effectively mitigate factual hallucinations in LLMs and improve the robustness of RALMs. However, there is still a risk that our method might generate hallucinations.
Appendix A Appendix
Models End-to-End External module Train data for LLM
(train) (inference) (data)
Self-Reasoning (Ours) Y N N No Need 2K
Self-RAG Asai et al. (2024) N N N No Need 145K (Generator)/ 46K (Critic)
ReAct Yao et al. (2023) N N Y No Need No Need
RECOMP Xu et al. (2023a) N Y Y 152K No Need
Table 3: Categorical comparisons with strong baseline models. External module (train) and External module (data) refer to whether the external module needs to be trained and the number of samples required, respectively. External module (inference) indicates whether the external module is needed during the inference stage. Train data for LLM indicates the number of training samples needed to train with LLMs.
A.1 More Related Work of LMs for Reasoning
One of the most well-known methods of using LLMs for reasoning is the Chain-of-Thought (CoT) Wei et al. (2022), which demonstrates the capability of LLMs to create their thinking process for problem-solving. Zhou et al. (2022) proposes a least-to-most prompting for solving complex tasks. Wang et al. (2022) introduces a method to reason with self-consistency. Press et al. (2023) proposes a method to further improve the chain of thought by reasoning explicitly instead of implicitly.
Recent works have extended beyond the internal reasoning ability of LLMs to include interactions with external tools (e.g., search engines or retrievers) for solving complex tasks. The ReAct Yao et al. (2023) presents an iterative paradigm to combine reasoning and acting with LLMs for tackling language reasoning and decision-making tasks. Xu et al. (2023b) introduces a framework to enable information retrieval and LLMs to interact with each other effectively with chain-of-query decomposition. Pan et al. (2024) proposes a novel framework named Chain-of-Action (CoA), which integrates a reasoning retrieval method to decompose complex questions into chains of configurable actions.
Different from the above works, which are mostly based on relatively large LLMs (e.g., ChatGPT), our proposed method focuses on enhancing smaller LLMs (e.g., LLaMA2) using only a limited number of samples to achieve high robustness and interpretability through single-step interaction.
A.2 Instructions
The instructions for GPT-4 to generate self-reasoning trajectories are shown in Figure 5 (the short-form and long-form QA tasks) and Figure 6 (the fact verification task). The words in the orange font are key fields that need to be generated.
A.3 Datasets Description
We conducted an extensive experimental evaluation of two short-form QA datasets, one long-form QA dataset, and a fact verification dataset.
NaturalQuestion (NQ) Kwiatkowski et al. (2019) contains real user questions issued to the Google search and answers found from Wikipedia by the annotators. NQ is created to train and evaluate automated question answering systems.
PopQA Mallen et al. (2023) is a large-scale open-domain question answering dataset, consisting of entity-centric QA pairs. Each question is made by converting a knowledge triplet retrieved from Wikidata using a template. In this work, we use PopQA to evaluate performance in long-tail settings.
ASQA Stelmakh et al. (2022) is a long-form factoid dataset, and most questions can be answered by Wikipedia. Each question originates from AmbigQA Min et al. (2020) and represents an ambiguous query that requires multiple short answers to cover various aspects. The dataset provides a long-form answer that contains all short answers.
FEVER Thorne et al. (2018) is a fact verification dataset that contains claims generated by rewriting sentences extracted from Wikipedia and subsequently verified without knowledge of the sentence from which they were derived. The claims are classified as Supported, Refuted, or NotEnoughInfo.
A.4 Experiment settings
Training settings.
During gradual learning, we fine-tune the LLaMA-2 Touvron et al. (2023) model with our self-reasoning framework for 3 epochs with a batch size set to 32, leveraging the DeepSpeed library Rasley et al. (2020) and the ZeRO optimizer, and we use parameter partitioning ZeRO stage 3 with float16 precision. The learning rate $r_{a}$ for the first stage is set to 5e-5, the learning rate $r_{b}$ for the second stage is set to 3e-5, and the learning rate $r_{c}$ for the final stage is set to 1e-5. Our self-reasoning 13B model is trained on the NVIDIA Tesla 8 $\times$ V100 32GB GPU for 4 hours, while the 7B model is trained for 2 hours.
Inference settings.
We use the vLLM  [^2] Codes are available at https://github.com/vllm-project/vllm framework Kwon et al. (2023) to accelerate the inference speed during inference. The codes follow the Apache-2.0 license agreement. We use greedy decoding in all experiments to ensure deterministic generations. We test the temperature within a range of $\{0.2,0.4,0.6,0.8,1.0\}$, finally we set the temperature to 0.2, as we observed lower temperature results in better performance in the open-domain question answering task. The maximum generation length is set to 2048 for our model. All baseline models are tested with zero-shot settings for short-form QA datasets, and with one-shot settings for the long-form QA and fact verification datasets.
Other settings.
For the document retrieval, we retrieve the top-$k$ relevant documents, and the $k$ is set to 5. We use the DPR and the Contriever in short-form QA settings. For long-form QA, we use GTR as a retrieval and evaluate it using one-shot to instruct the model to generate citations. For the data generation quality control setting, the threshold for citation recall is set to 0.8, and the threshold of citation precision is set to 0.8.
A.5 Categorical Comparisons
We differentiate our method from existing strong baseline models by categorizing and comparing it across six dimensions, as presented in Table 3. As illustrated in the table, our method can stand out in several key aspects.
First, our self-reasoning method is the only end-to-end framework among existing methods that can improve performance without relying on external models or tools. Second, our method eliminates the need for external modules during both the training and inference phases. In practical applications, our framework does not need to call multiple tools or modules. Third, our framework requires a significantly smaller dataset for training the LLM compared to other methods, needing only 2,000 samples with self-reasoning trajectories. This efficiency in training drastically lowers the resources and time needed, making our method both cost-effective and scalable for practical applications.
A.6 Case Study
In our case study, as illustrated in Figure 7, we compare the responses generated by the raw LLM, the standard RALM (e.g. LLaMA2 with retrievals), and our self-reasoning method. The challenge involves reconciling information from multiple retrieved documents to provide a correct answer as the retrieved documents contained noisy data.
The response from the raw LLM (e.g., LLaMA2) suggested that the film was made in 2000, based on its inherited knowledge. However, this answer is incorrect and is a hallucination generated by the LLM. The standard RALM approach yielded 1989 as the production date. This answer was based on unrelated details from the retrieved documents, showing a lack of context-specific understanding and robustness for noisy retrieved documents.
Our self-reasoning framework provided a comprehensive approach by assessing the relevance and context of retrieved documents. First, in the relevance-aware process, the documents were identified as relevant based on their content regarding the production dates and events surrounding the film. Second, in the evidence-aware selective process, the model retrieved the first documents, which highlighted the original start date as January 2002, with filming commencing in February 2002 (highlighted in green in the figure). This information was crucial in establishing the timeline for the film’s production. The model can also understand of the difference between the production date and the release date in the third retrieved document (highlighted in red in the figure). In the trajectory analysis process, the correct timeline was deduced by piecing together self-generated trajectories, leading to the conclusion that the film Catch Me If You Can was indeed produced in 2002. As the case illustrated above, by leveraging relevant documents and focusing on contextual evidence, our self-reasoning framework can achieve a precise and well-supported answer, highlighting its utility and robustness in complex information retrieval tasks.
A.7 More Analysis on Ablation Study
Effectiveness of Gradual Learning
Further, we validate the effect of gradual learning. Rather than training an LLM with a stage-by-stage approach, we initially concatenate the reasoning trajectories from all stages and put them into the LLM for end-to-end training. As shown in Table 2, the performance decline can be observed in three datasets, suggesting that gradual learning can help improve performance.
Effectiveness of Quality Control
The effect of quality control on data generation is also evaluated in our work. Instead of using the filtered high-quality training samples, we randomly sampled 2,000 unfiltered training samples generated by GPT-4. As shown in Table 2, the substitution of unfiltered training data leads to the degradation of the model results.
A.8 Human Evaluation
We randomly sample 100 examples from the ASQA dataset and annotate the outputs of selected models. Each sample is then assigned to two people for annotation. Each annotator is required to verify the citation recall and citation precision acorrding to our provided scheme. The annotation scheme is inspired by Gao et al. (2023b) as follows:
Citation Recall.
The annotators are shown the question $q$, the statement $s_{i}$, and all of its citations $C_{i}$, and they assess if the set of citations fully support the statement (recall=1) or if they do not support all the claims (recall=0). We calculate the overall recall score for the model by averaging the recall scores of all statements.
Citation Precision
The annotators are shown the question $q$ and a statement $s_{i}$ and one of its citation ${c^{(k)}_{i}}\in C_{i}$. We ask the annotator if the citation fully supports, partially supports, or does not support the generated claims in $s_{i}$. Citation ${c^{(k)}_{i}}$ has a citation precision of 1 if $s_{i}$ has a recall of 1, and ${c^{(k)}_{i}}$ fully or partially supports $s_{i}$. Finally, we calculate the overall precision score for the model by averaging the precision scores of all statements.
Figure 5: The instructions for the GPT-4 to generate the self-reasoning trajectories for short-form and long-form QA tasks.
Figure 6: The instructions for the GPT-4 to generate the self-reasoning trajectories for the fact verification task.
Figure 7: A Case Study. We present the self-reasoning trajectories generated by our framework during inference and demonstrate how they can logically generate the correct answer.

### v3

@@
 Abstract
-The Retrieval-Augmented Language Model (RALM) has shown remarkable performance on knowledge-intensive tasks by incorporating external knowledge during inference, which mitigates the factual hallucinations inherited in large language models (LLMs).
-Despite these advancements, challenges persist in the implementation of RALMs, particularly concerning their reliability and traceability.
-To be specific, the irrelevant document retrieval may result in unhelpful response generation or even deteriorate the performance of LLMs, while the lack of proper citations in generated outputs complicates efforts to verify the trustworthiness of the models.
+The Retrieval-Augmented Language Model (RALM) has demonstrated remarkable performance on knowledge-intensive tasks by integrating external knowledge during inference, which mitigates the factual hallucinations inherited in large language models (LLMs).
+Despite these advancements, challenges persist in the implementation of RALMs, particularly in terms of reliability and traceability.
+Specifically, the irrelevant document retrieval may result in unhelpful responses or even deteriorate the performance of LLMs, while the lack of appropriate citations in outputs complicates efforts to verify the trustworthiness of the models.
 To this end, we propose a novel self-reasoning framework aimed at improving the reliability and traceability of RALMs, whose core idea is to leverage reasoning trajectories generated by the LLM itself.
-The framework involves constructing self-reason trajectories with three processes: a relevance-aware process, an evidence-aware selective process, and a trajectory analysis process.
-We have evaluated our framework across four public datasets (two short-form QA datasets, one long-form QA dataset, and one fact verification dataset) to demonstrate the superiority of our method, which can outperform existing state-of-art models and can achieve comparable performance with GPT-4, while only using 2,000 training samples.
-Improving Retrieval Augmented Language Model with Self-Reasoning
-Yuan Xia, Jingbo Zhou, Zhenhui Shi, Jun Chen, Haifeng Huang Baidu Inc., China. {xiayuan,zhoujingbo,shizhenhui,chenjun22,huanghaifeng}@baidu.com
-1 Introduction
-Figure 1: An example of how self-reasoning framework generates reasoning trajectories with a relevant-aware process, an evidence-aware selective process, and a trajectory analysis process.
-The Retrieval-Augmented Language Model (RALM) has become a critical enhancement for Large Language Models (LLMs) by integrating external knowledge during inference.
-Despite their advanced capabilities in language understanding and generation Brown et al.
-(2020); Touvron et al.
-(2023), LLMs are prone to producing hallucinated and inaccurate content, particularly in knowledge-intensive tasks Ji et al.
-(2023).
-Augmenting LLMs with relevant information obtained from external sources like Wikipedia and search engines has proven effective in reducing these inaccuracies Guu et al.
-(2020); Lewis et al.
-(2020); Borgeaud et al.
-(2022); Izacard et al.
-(2022); Asai et al.
-(2024).
-This method has proven effective in mitigating the factual hallucinations inherent in LLMs Kwiatkowski et al.
-(2019); Petroni et al.
-(2021); Ram et al.
-(2023).
-Nevertheless, there are still limitations associated with RALMs, particularly in terms of reliability and traceability.
-First, the reliability of the retrieved information remains a significant concern.
-Previous studies have shown that noisy retrieval can adversely affect the performance of an LLM Menick et al.
-(2022); Li et al.
-(2023), as irrelevant data can lead to misguided responses and disturb the model’s ability to leverage its intrinsic knowledge effectively.
-Second, the interpretability and traceability of outputs generated by RALMs need to be improved.
-While RALMs incorporate retrieved documents during both training and inference phases, they may fail to cite these documents explicitly,
-thus complicating the process of tracing and verifying the claims made by LLMs.
-To improve the retrieval robustness, recent studies have explored incorporating natural language inference (NLI) models Honovich et al.
-(2022) and document summarization models during inference Yoran et al.
-(2023); Xu et al.
-(2023a).
-However, the effectiveness of these external NLI and summarization models largely affects the overall performance of RALMs.
-The training and optimization of these auxiliary models require additional costs.
-Consequently, determining the most appropriate training and selection methods for these NLI and summarization models presents a further challenge in the application of such approaches.
+The framework involves constructing self-reasoning trajectories through three processes: a relevance-aware process, an evidence-aware selective process, and a trajectory analysis process.
+We evaluated our framework across four public datasets (two short-form QA datasets, one long-form QA dataset, and one fact verification dataset) to demonstrate its superiority.
+Our method can outperform existing state-of-the-art models and achieve performance comparable with GPT-4, using only 2,000 training samples.
+Introduction
+Figure 1: An example of how self-reasoning framework generates reasoning trajectories.
+The Retrieval-Augmented Language Model (RALM), also known as Retrieval-Augmented Generation (RAG), has become a crucial enhancement for Large Language Models (LLMs) by integrating external knowledge during inference.
+Despite their advanced capabilities in language understanding and generation (Brown et al.
+2020; Touvron et al.
+2023), LLMs are prone to producing hallucinated and inaccurate content, especially in knowledge-intensive tasks (Ji et al.
+2023).
+Augmenting LLMs with relevant information obtained from external sources like Wikipedia and search engines has proven effective in reducing these inaccuracies (Guu et al.
+2020; Lewis et al.
+2020; Borgeaud et al.
+2022; Izacard et al.
+2022; Asai et al.
+2024).
+This approach has proven effective in mitigating the factual hallucinations that are inherent in LLMs (Kwiatkowski et al.
+2019; Petroni et al.
+2021; Ram et al.
+2023).
+Nevertheless, there are still limitations associated with RALMs, particularly concerning reliability and traceability.
+Firstly, the reliability of the retrieved information remains a substantial concern.
+Previous studies have shown that noisy retrieval can adversely affect the performance of an LLM (Menick et al.
+2022; Li et al.
+2023), as irrelevant data can lead to misguided responses and disturb the model’s ability to leverage its intrinsic knowledge effectively.
+Secondly, the interpretability and traceability of outputs generated by RALMs need to be improved.
+Although RALMs incorporate retrieved documents during both the training and inference phases, they may fail to explicitly cite these documents, thus complicating the process of tracing and verifying the claims made by LLMs.
+To improve the retrieval robustness, recent studies have explored incorporating external tools such as natural language inference (NLI) models (Honovich et al.
+2022) and document summarization models during inference (Yoran et al.
+2023; Xu et al.
+2024).
+However, the effectiveness of these external tools largely influences the overall performance of RALMs.
+Additionally, training and optimizing these auxiliary models require additional costs.
+Consequently, identifying the most appropriate training and selection methods for NLI and summarization models remains a critical challenge in leveraging these approaches.
 To address the above limitations, we propose a novel end-to-end self-reasoning framework to improve the performance of RALMs.
-Our intuition is that the explicit self-reasoning trajectory crafted by LLMs can improve the retrieval robustness and accuracy of question answering.
-During the pre-training phase, an LLM primarily focuses on knowledge acquisition, yet it does not learn to reason from retrieved documents to generate answers.
+For convenience, we will also refer to this framework as self-reasoning RAG and use the terms interchangeably.
+Our intuition is that the explicit self-reasoning trajectory crafted by LLMs can improve both the retrieval robustness and accuracy in question answering.
+During the pre-training phase, while an LLM primarily focuses on knowledge acquisition, it does not learn to reason from retrieved documents to generate answers.
 To address this, a feasible approach is to incorporate reasoning trajectories into a post-training phase.
-Such an approach could potentially teach the model to reason and distinguish relevant and irrelevant documents, improving its response accuracy to queries.
+Such an approach could potentially teach the model to reason and distinguish relevant and irrelevant documents, thereby enhancing its query response accuracy.
 An example of how our self-reasoning framework generates reasoning trajectories is illustrated in Figure 1.
 In contrast, as shown in the middle part of Figure 2, the conventional RALM methods gather all documents in a non-selective manner, leading to the distraction of the LLM by irrelevant content and consequently resulting in the generation of erroneous answers.
-Our framework comprises three self-reasoning processes: 1) a Relevance-Aware Process (RAP), which instructs the LLM to judge the relevance between the retrieved documents and the question, 2) an Evidence-Aware Selective Process (EAP), which directs the LLM to choose and cite relevant documents, and then automatically select snippets of key sentences as evidence from the cited documents, 3) a Trajectory Analysis Process (TAP), which requires the LLM to generate a concise analysis based on all gathered self-reasoning trajectories generated by previous two processes and subsequently provide the final inferred answer.
+Our framework constructs self-reasoning trajectories comprising three processes: 1) a Relevance-Aware Process (RAP), which instructs the LLM to judge the relevance between the retrieved documents and the question, 2) an Evidence-Aware Selective Process (EAP), which directs the LLM to choose and cite relevant documents, and then automatically select snippets of key sentences as evidence from the cited documents, 3) a Trajectory Analysis Process (TAP), which requires the LLM to synthesize a concise analysis based on all gathered self-reasoning trajectories generated by previous two processes and subsequently provide the final inferred answer.
 Furthermore, we propose a gradual training method by employing stage-wise masking strategies to enhance the performance of our framework.
 We summarize our contributions as follows:
-  • We propose a novel end-to-end self-reasoning framework that improves the robustness of RALMs by leveraging reasoning trajectories generated by the LLM itself, without the need for external models or tools.
+  • We propose a novel end-to-end self-reasoning framework that improves the robustness of RALMs by leveraging reasoning trajectories generated by the LLM itself, without the need for external tools.
   • We carefully design three processes to enhance the interpretability and traceability of RALMs by requiring LLMs to explicitly generate snippets and citations from documents, and further explain the reason why cited documents can help answer the question.
-  • We evaluate our framework on four public datasets (two short-form QA, one long-form QA, and one fact verification), demonstrating that our method surpasses existing state-of-art models in performance, achieving this with only using 2,000 training samples.
-Figure 2: An illustration of the self-reasoning framework for improving the RALMs.
+  • We evaluate our framework on four public datasets (two short-form QA, one long-form QA, and one fact verification), demonstrating that our method surpasses existing state-of-the-art models in performance using only 2,000 training samples.
+Figure 2: An illustration of the self-reasoning framework.
 The upper is the basic LLMs which answer the question by inherent knowledge.
@@
 The bottom is our self-reasoning framework which uses self-generated reason trajectories to output answers.
-2 Related Work
-2.1 Retrieval-augmented LMs
-Many studies have investigated augmenting the performance of LLMs with externally retrieved information Izacard et al.
-(2022); Guu et al.
-(2020); Borgeaud et al.
-(2022) and some of them pre-train language models with retrieved passages.
+Related Work
+Retrieval-augmented LMs
+Many studies have investigated augmenting the performance of LLMs with externally retrieved information (Izacard et al.
+2022; Guu et al.
+2020; Borgeaud et al.
+2022) and some of them pre-train language models with retrieved passages.
 For works focusing on RALMs with citations, Menick et al.
@@
 (2023b) proposes an end-to-end system to retrieve supporting evidence and generate answers with citations, while only focusing on prompting without updating their model weights.
-Other works instruct or fine-tune LLMs to use external tools to retrieve dynamically Schick et al.
-(2023); Yao et al.
-(2023); Jiang et al.
-(2023), which offers an adaptive method of when and what to search.
+Other works instruct or fine-tune LLMs to use external tools to retrieve dynamically (Schick et al.
+2023; Yao et al.
+2023; Jiang et al.
+2023), which offers an adaptive method of when and what to search.
 Gao et al.
 (2023a) improves the attribution and factuality of language models by taking outputs of LLMs and applying a post-process retrieve-and-edit approach.
-2.2 Robustness for RALMs
-To improve the robustness of RALMs, Yoran et al.
-(2023) use a natural language inference model to filter out irrelevant documents before RALMs, and Xu et al.
-(2023a) use a memorization model to filter out or compress retrieved documents before using them to prompt an LLM.
-The concurrent work Yu et al.
-(2023) generates summaries of documents as notes before RALMs, and Baek et al.
-(2023) uses a separate small language model as a verifier to detect and correct errors in LLMs during retrieval for generating factually correct outputs.
-Different from the above works, ours identifies key sentences and cites relevant documents through an end-to-end framework that eliminates irrelevant ones, without dependence on external inference models.
-The most similar work to ours is presented by Asai et al.
-(2024), who developed a method that teaches models to retrieve information using designed reflection tokens.
+Robustness for RALMs
+To improve the robustness of RALMs, previous works can be divided into two categories.
+The first category utilizes retrieved documents to enhance the Chain of Thought (CoT).
+For example, IRCoT (Trivedi et al.
+2023) iteratively uses retrieved documents to generate CoT, which is then used to retrieve further documents in subsequent steps.
+ReAct (Yao et al.
+2023) introduces an iterative CoT paradigm that integrates reasoning with search results.
+However, irrelevant retrievals may produce misguided CoT, adversely affecting LLM performance (Menick et al.
+2022; Li et al.
+2023).
+To address the issue of irrelevant retrieval information, the second category proposes using external modules to process retrieved documents during inference.
+For instance, Yoran et al.
+(2023) utilize a natural language inference model to filter out irrelevant documents, Yan et al.
+(2024) employ a retrieval evaluator to classify documents based on their quality, and Xu et al.
+(2024) and Yu et al.
+(2023) apply models to filter out or compress retrieved documents.
+Baek et al.
+(2023) deploy a separate small language model as a verifier to detect and correct errors in LLMs during retrieval.
+A method presented by Asai et al.
+(2024), which appears most similar to our approach, develops a technique that instructs models to retrieve information using specifically designed reflection tokens.
 However, this approach needs to train extra critic models and generator models to predict the reflection tokens, which requires tens of thousands of extra training samples.
-In contrast, our method does not rely on special tokens.
-Instead, we construct reasoning trajectories and then use them to directly enhance the performance of LLMs, offering a more efficient and scalable solution.
-More related works on LLMs for reasoning are discussed in the Appendix §$\S$§A.1.
-3 Preliminary
+Unlike the second group of works, which rely on external tools or additional modules to eliminate irrelevant information, the self-reasoning RAG method integrates self-reasoning directly into the model’s architecture, thereby enhancing the performance of LLMs and providing a more efficient and scalable solution.
+Further related works on LLMs for reasoning are discussed in the Appendix.
+Preliminary
 We formally define the problem of retrieval augmented generation with self-reasoning.
@@
 In our work, we train an LLM (e.g.
-LLaMA2) to first generate reasoning trajectories $\tau$ through self-reasoning and then to generate answers ${y}^{*}$ (including long-form answers and short-form answers) on condition of $\tau$.
+LLaMA2) to first generate reasoning trajectories $\tau$ through self-reasoning and then to generate answers ${y}^{*}$ (short-form answers) on condition of $\tau$.
 The model output is $y={\rm{concat}}(\tau,y^{*})$, which is the concatenation of $\tau$ and $y^{*}$.
-Note that the generations of $\tau$ and $y^{*}$ are done in a single pass within our self-reasoning framework.
-4 Method
-In this section, we provide a detailed implementation of the self-reasoning process and the method to construct reasoning trajectories.
-In this work, we propose a novel framework which involves three processes: 1) a Relevance-Aware Process (Sec §$\S$4.1), 2) an Evidence-Aware Selective Process (Sec §$\S$4.2), and 3) a Trajectory Analysis Process (Sec §$\S$4.3).
-An illustration of our self-reasoning framework for improving the RALMs is shown in Figure 2.
-In Sec 4.4, we demonstrate the process of data generation and quality control, and Sec 4.5 presents the details of model training.
-4.1 Relevance-Aware Process
-In this work, we choose the DPR Karpukhin et al.
-(2020) and the Contriever Izacard et al.
-(2021) as the default retriever $R$ to recall the top-$k$ relevant documents.
+Note that the generations of $\tau$ and $y^{*}$ are done in a single pass within the self-reasoning framework.
+Method
+Here we provide a detailed implementation of the self-reasoning process which involves three processes: 1) a Relevance-Aware Process (RAP), 2) an Evidence-Aware Selective Process (EAP), and 3) a Trajectory Analysis Process (TAP).
+An illustration of our self-reasoning framework is shown in Figure 2.
+Additionally, we outline the process of data generation and quality control, and present the specifics of model training.
+Relevance-Aware Process
+In this work, we choose DPR (Karpukhin et al.
+2020) and Contriever (Izacard et al.
+2021) as default retrievers $R$ to recall the top-$k$ relevant documents.
 When presented with a question and a set of documents, people can determine whether the question is relevant to the retrieved documents.
@@
 The output should include two fields as relevant and relevant reason, as depicted in Figure 2.
-Noting that if all of the retrieved documents are irrelevant, the model should provide an answer based on the internal knowledge acquired during its pre-training phase.
+If all of the retrieved documents are irrelevant, the model should provide an answer based on the internal knowledge acquired during its pre-training phase.
 We define the self-reasoning trajectories generated by RAP as $\tau_{r}$.
-4.2 Evidence-Aware Selective Process
-When required to answer the question, people generally will first identify the crucial sentences from the provided documents, and then cite or highlight them as key points.
+Evidence-Aware Selective Process
+When answering a question, people generally first identify the crucial sentences from the provided documents and then cite or highlight them as key points.
 This process of citing the document facilitates reading comprehension and can serve as a technique for combining multiple short answers to address various aspects.
@@
 We define the self-reasoning trajectories generated by EAP as $\tau_{e}$.
-4.3 Trajectory Analysis Process
+Trajectory Analysis Process
 Finally, we consolidate all the self-reasoning trajectories ($\tau_{r}$ and $\tau_{e}$) in the previous processes together to form a chain of reasoning snippets, thereby enhancing the overall performance of the retrieval augmentation generation.
@@
 We define the self-reasoning trajectories generated by TAP as $\tau_{a}$.
-In this work, the analysis output is defined as a long-form answer, and the answer output is defined as a short-short answer.
-In Section 5.2, we further explored the performance of long-form and short-form QA settings.
-4.4 Data Generation and Quality Control
+In this work, the analysis output is defined as a long-form answer, and the answer output is defined as a short-form answer.
+In the experiment section, we further explored the performance of long-form and short-form QA settings.
+Data Generation and Quality Control
 Training Data Generation.
-For the Relevance-Aware Process data generation, as manually labeling the relevant and irrelevant documents is label-intensive, we request the GPT-4 OpenAI (2023) to generate answers as ground truth.
-Specifically, we instruct the GPT-4 to generate labels regarding irrelevant fields, and further to output the reasons why the given documents cannot answer the question.
+For the Relevance-Aware Process data generation, as manually labeling the relevant and irrelevant documents is label-intensive, we request GPT-4 (OpenAI 2023) to generate answers as ground truth.
+Specifically, we instruct GPT-4 to generate labels regarding irrelevant fields, and further to output the reasons why the given documents cannot answer the question.
 We concatenate the given question and the retrieved documents as positive samples.
@@
 To avoid order bias in the training data, we shuffle the order of the documents.
-For the EAP and TAP data generation, manually annotating the citation of the document and writing the self-reasoning process for each question is not feasible in practice.
-Therefore, we follow a similar process to RAP, we first instruct the GPT-4 to generate a snippet of selected documents and subsequently output the reasoning process as trajectories.
-The method to construct the EAP training data is the same as RAP except that the instructions for the GPT-4 are different.
-The details of the instructions are shown in Appendix §$\S$§A.2.
+For the EAP and TAP data generation, manually annotating the citation and writing the self-reasoning process for each question is not feasible in practice.
+Therefore, we follow a similar process to RAP, we first instruct GPT-4 to generate a snippet of selected documents and subsequently output the reasoning process as trajectories.
+The method for constructing the EAP training data is the same as RAP except that the instructions given to GPT-4 are different.
+The details of the instructions are shown in the Appendix.
 Data Quality Control.
 For training data generation, correct and comprehensive reasoning trajectories are very important.
-When training an LLM, the quality of the training samples is more important than the quantity Zhou et al.
-(2023).
-As we cannot guarantee the correctness of self-reasoning trajectories and citations by the GPT-4, we develop two efficient methods to control the quality of data generation: 1) The first method is to use the off-the-shelf tools  [^1] Tools are available at https://github.com/princeton-nlp/ALCE/tree/main in Gao et al.
+When training an LLM, the quality of the training samples is more important than the quantity (Zhou et al.
+2023).
+As we cannot guarantee the correctness of self-reasoning trajectories and citations by GPT-4, we develop two efficient methods to control the quality of data generation: 1) The first method is to use the off-the-shelf tools  [^1] Tools are available at https://github.com/princeton-nlp/ALCE/tree/main in Gao et al.
 (2023b) to automatically verify the performance of data generation for document citations.
@@
 We totally generate 10,000 training samples by GPT-4, after the filtering strategy by quality control, we finally keep 2,000 training samples with high quality.
-4.5 Model Training
-We train our self-reasoning generation model $\phi$ by our constructed corpus which is augmented with self-reasoning trajectories $\tau$ using the standard language modeling objective, maximizing likelihood:
+More details and pseudo-codes can be found in the Appendix.
+Model Training
+We train the self-reasoning RAG model $\phi$ by our constructed corpus which is augmented with self-reasoning trajectories $\tau$ using the standard language modeling objective, maximizing likelihood:
 $\max_{\mathcal{\phi}}\mathbb{E}_{(q,\tau,y)\sim\mathcal{D}_{sr}}\log p_{\mathcal{\phi}}(y\mid\tau,q)p_{\mathcal{\phi}}({\tau\mid q})$ (1)
-where $\tau=\tau_{r}\oplus\tau_{e}\oplus\tau_{a}$ are the self-reasoning trajectories, $\oplus$ is the concatenation operator, $\tau_{r},\tau_{e},\tau_{a}$ are trajectories generated by above three processes respectively.
+where $\tau=\tau_{r}\oplus\tau_{e}\oplus\tau_{a}$ are the self-reasoning trajectories, $\oplus$ is a concatenation operator, $\tau_{r},\tau_{e},\tau_{a}$ are trajectories generated by above three processes respectively.
 $q$ is the provided question, and $y$ is the model output, including the intermediate reason trajectories and the final answer.
@@
 Finally, we concatenate the reasoning trajectories from all stages and put them into a self-reasoning LLM for end-to-end training with a learning rate $r_{c}$.
-Hyper-parameters for training are described in Appendix §$\S$§A.4.
-5 Experiments
+Hyper-parameters for training are described in the Appendix.
+Experiments
 Models NaturalQuestion PopQA FEVER ASQA
@@
 LLaMA213B-chat 32.7 53.5 53.4 26.4 39.4 38.4
-Vicuna7B Chiang et al. (2023) 28.0 55.2 62.4 24.3 45.7 40.8
-Vicuna13B Chiang et al. (2023) 35.4 56.1 60.6 27.3 51.3 50.2
+Vicuna7B (Chiang et al. 2023) 28.0 55.2 62.4 24.3 45.7 40.8
+Vicuna13B (Chiang et al. 2023) 35.4 56.1 60.6 27.3 51.3 50.2
 LLaMA2-FT7B 36.8 54.4 67.5 28.5 47.2 45.4
-ReAct Yao et al. (2023) - - 64.6 - - -
-RECOMP Xu et al. (2023a) 38.4 - - - - -
-Self-RAG7B Asai et al. (2024) 37.2 54.9 70.2 30.0 66.9 67.8
-Self-RAG13B Asai et al. (2024) 38.8 55.8 72.1 31.7 70.3 71.3
+ReAct (Yao et al. 2023) - - 64.6 - - -
+RECOMP (Xu et al. 2024) 38.4 - - - - -
+Self-RAG7B (Asai et al. 2024) 37.2 54.9 70.2 30.0 66.9 67.8
+Self-RAG13B (Asai et al. 2024) 38.8 55.8 72.1 31.7 70.3 71.3
 Self-Reasoning7B 38.0 54.2 78.6 33.9 66.3 70.8
@@
 The results are averaged over five runs, and presented with standard variance values omitted (all $\leq 2\%$).
-5.1 Datasets and Settings
-To demonstrate the effectiveness of our proposed self-reasoning framework, we conduct an extensive experimental evaluation on two short-form QA datasets (NaturalQuestion Kwiatkowski et al.
-(2019) and PopQA Mallen et al.
-(2023)), one long-form QA dataset (ASQA Stelmakh et al.
-(2022)), and one fact verification dataset (FEVER Thorne et al.
-(2018)).
-Detailed descriptions of the datasets can be found in Appendix §$\S$§A.3.
+Datasets and Settings
+To demonstrate the effectiveness of our proposed self-reasoning framework, we conduct an extensive experimental evaluation on two short-form QA datasets (NaturalQuestion (Kwiatkowski et al.
+2019) and PopQA (Mallen et al.
+2023)), one long-form QA dataset (ASQA (Stelmakh et al.
+2022)), and one fact verification dataset (FEVER (Thorne et al.
+2018)).
+Detailed descriptions of the datasets can be found in the Appendix.
 We explore off-the-shelf retrievers.
-We use the DPR Karpukhin et al.
-(2020) and the Contriever-MS MARCO Izacard et al.
-(2021) to retrieve the top five documents from Wikipedia.
-By default, we use the DPR as a retriever for the NQ, as the DPR has been fine-tuned on the high-quality NQ data.
-On the PopQA, where question and answer pairs are created based on Wikipedia in 2022, therefore, for the PopQA, we use the December 2020 preprocessed Wikipedia corpus provided by Izacard et al.
-(2022) and use the Contriever as a retriever.
-For the ASQA dataset, we use GTR Ni et al.
-(2022) as a retrieval that corresponds to the experimental settings in Gao et al.
-(2023b).
-More settigns can be found in Appendix §$\S$§A.4.
-5.2 Evaluation Metrics
+We use DPR (Karpukhin et al.
+2020) and Contriever-MS MARCO (Izacard et al.
+2021) to retrieve the top five documents from Wikipedia.
+By default, we use DPR as a retriever for the NQ, as DPR has been fine-tuned on the high-quality NQ data.
+On the PopQA, where question and answer pairs are created based on Wikipedia in 2022, therefore, for the PopQA, we use the December 2020 preprocessed Wikipedia corpus provided by (Izacard et al.
+2022) and use Contriever as a retriever.
+For the ASQA dataset, we use GTR (Ni et al.
+2022) as a retrieval that corresponds to the experimental settings in (Gao et al.
+2023b).
+More settings can be found in the Appendix.
+Evaluation Metrics
 We use different evaluation metrics for short-form QA, long-form QA, and fact verification tasks.
@@
 Long-form QA metrics.
-For long-form QA tasks, we report the EM recall as a correctness metric, and the citation recall and the citation precision for citation quality, which are the same as the metrics in Gao et al.
-(2023b).
+For long-form QA tasks, we report the EM recall as a correctness metric, and the citation recall and the citation precision for citation quality, which are the same as the metrics in (Gao et al.
+2023b).
 Fact verification metrics.
 For the fact verification task, we report the accuracy as a metric, which is a three-class classification accuracy, following Thorne et al. (2018).
-5.3 Baseline Models
+Baseline Models
 Baseline models without retrieval.
 We evaluate strong open-source pre-trained LLMs as baseline models.
-For basic LLMs, we test LLaMA2-7B, LLaMA2-13B Touvron et al.
-(2023) and its instruction-tuned chat version LLaMA2-Chat-7B, LLaMA2-Chat-13B.
+For basic LLMs, we test LLaMA2-7B, LLaMA2-13B (Touvron et al.
+2023) and its instruction-tuned chat version LLaMA2-Chat-7B, LLaMA2-Chat-13B.
 Baseline models with retrieval.
-To assess the performance of the retrieval augmented LMs, we use retrievers described in Section 5.1.
-First, we benchmark the models using the LLaMA2 and the Vicuna Chiang et al.
-(2023) series models for baselines.
+First, we benchmark the models using the LLaMA2 and the Vicuna (Chiang et al.
+2023) series models for baselines.
 Additionally, for a fair comparison, we also include LLaMA2-FT, where LLaMA2 is fine-tuned on all the training samples generated by GPT-4 except the self-reasoning trajectories.
-To establish strong baselines, we compare our method against RECOMP Xu et al.
-(2023a), ReAct Yao et al.
-(2023), and Self-RAG Asai et al.
-(2024), all of which are trained with extra GPT-4 generated samples or external tools.
-We also compare our framework with GPT-4 OpenAI (2023).
-We include categorical comparisons with the baseline models in the Appendix §$\S$§A.5.
-Figure 3: Noise robustness experiment results on three different datasets: (a) On the left is the NQ dataset, (b) in the middle is the PopQA dataset, (c) on the right is the FEVER dataset.
+To establish strong baselines, we compare our method against RECOMP (Xu et al.
+2024), ReAct (Yao et al.
+2023), and Self-RAG (Asai et al.
+2024), all of which are trained with extra GPT-4 generated samples or external tools.
+We also compare our framework with GPT-4 (OpenAI 2023).
+We include categorical comparisons with the baseline models in the Appendix.
+Figure 3: Noise robustness experiment results on three different datasets: (a) On the left is the NQ dataset, (b) in the middle is the PopQA dataset, and (c) on the right is the FEVER dataset.
 The Self-RAG and Vicuna are 13B parameter size models.
-5.4 Main Results
+Main Results
 Table 1 shows the performance comparisons with different methods on the four public datasets.
@@
 The EAP and TAP are specifically designed for multi-document reading comprehension, enabling our performance to surpass other baselines.
-In terms of citation evaluation metrics, our self-reasoning framework can achieve better results than GPT-4 in ASQA citation recall metrics (72.3 vs.
+In terms of citation evaluation metrics, the self-reasoning RAG can achieve better results than GPT-4 in ASQA citation recall metrics (72.3 vs.
 68.5).
-This is largely due to the reasoning trajectories generated in the EAP, which can enhance the recall and precision of citation evaluation, leading to more interpretable generations.
+This is largely due to the reasoning trajectories generated in the EAP, which can enhance the recall and precision of citation evaluation, leading to more interpretable and traceable generations.
 For fact verification evaluations, we observed that self-seasoning is dominantly superior to all baseline models.
@@
 72.1).
-The RAP in our framework is designed to judge the relevance between the retrieved documents and the question, which leads to the notable enhancement in accuracy for this fact verification task.
-To clearly demonstrate the practical applications and benefits of our self-reasoning framework, we provide a case study for a more in-depth analysis in the Appendix §$\S$§A.6, which illustrates how the framework operates in real-world scenarios.
+The RAP in our framework is designed to judge the relevance between the retrieved documents and the question, which leads to a notable enhancement in accuracy for this fact verification task.
+To clearly demonstrate the practical applications and benefits of our self-reasoning framework, we provide a case study for a more in-depth analysis in Appendix, which illustrates how our framework operates in real-world scenarios.
 Models NQ PopQA FEVER
@@
 In the table, the origin represents our self-reasoning model enhanced with self-generated trajectories.
-6 Analysis
-6.1 Ablation Study
+Models NQ PopQA FEVER
+(acc) (acc) (acc)
+LLaMA2 32.7 53.5 53.4
++ trajectory 38.3 54.2 79.2
+[…]
++ trajectory 38.5 56.4 79.6
+Table 3: The analysis on the effectiveness of self-reasoning trajectories with 13B parameter size models.
+In the table, the +trajectory indicates the result of the baseline model is enhanced with self-generated trajectories by our framework.
+Analysis
+Ablation Study
 We conduct an ablation study on two short-form QA datasets and a fact verification dataset to analyze the individual contributions of each process within our proposed self-reasoning framework.
-We further explore the effectiveness of the gradual learning (GL) method and the quality control (QC) of data generation (a detailed analysis described in Appendix §$\S$§A.7).
-The main ablation study results are shown in Table 2.
-Effectiveness of Relevant-Aware Process
-First, we evaluate the effect of the Relevant-Aware Process (RAP).
+We further explore the effectiveness of the gradual learning (GL) method and the quality control (QC) of data generation (a detailed analysis described in the Appendix).
+The main ablation study results are shown in Table 2 and Table 3.
+Effectiveness of RAP.
+First, we evaluate the effect of the RAP.
 The removal of the RAP causes the overall performance to drop in two short-form QA datasets and a fact verification dataset, suggesting that preliminary consideration of the relevance between questions and retrieved documents can help improve performance.
@@
 Our model will immediately output NotEnoughInfo if it detects that all documents are irrelevant.
-Effectiveness of Evidence-Aware Process
-Then we evaluate the effect of the Evidence-Aware Selective Process (EAP).
-Removing the EAP causes the overall performance of the average accuracy to decline from 60.9 to 56.3 in three short-form QA datasets.
-This reduction indicates that snippets of key sentences and document citations generated through self-reasoning are instrumental in boosting accuracy.
-Effectiveness of Trajectory Analysis Process
-Finally, we evaluate the effect of the Trajectory Analysis Process (TAP).
+Effectiveness of EAP.
+Then we evaluate the effect of the EAP.
+Removing the EAP causes the overall performance of the average accuracy to decline from 60.9 to 56.3 in three short-form QA datasets, which indicates that snippets of key sentences and document citations generated through self-reasoning are instrumental in boosting accuracy.
+Effectiveness of TAP.
+Finally, we evaluate the effect of the TAP.
 When excluding the TAP, we can observe a performance decline on all three datasets, demonstrating that self-analysis based on two previous processes generated trajectories can also improve the performance of LLMs.
 Note that the analysis content generated by TAP is indispensable for the long-form QA evaluation.
-6.2 Retrieval Robustness Analysis
-As retrievers are not perfect and past work has shown that noisy retrieval can have negative effects on the performance of LLMs Petroni et al.
-(2020); Li et al.
-(2023).
+Effectiveness of Self-Reasoning Trajectory.
+To verify whether the trajectories generated by the self-reasoning framework are truly effective, we put the trajectories generated by our self-reasoning framework into the original baseline models as input prompts, and then use the baseline models to regenerate the answers.
+We observe that incorporating self-generated trajectories can significantly enhance performance in short QA tasks and fact verification tasks.
+Retrieval Robustness Analysis
+Retrievers are not perfect and past work has shown that noisy retrieval can have negative effects on the performance of LLMs (Petroni et al.
+2020; Li et al.
+2023).
 In this section, we design two kinds of settings to validate the robustness of RALMs.
 In the first setting, we test whether the order of the retrieved documents will affect the performance of the RALMs.
-Specifically, after retrieving the top-$k$ documents using retrievals (such as the DPR) with a descending relevance score, we randomly shuffle the order of the retrieved documents and then input them to an LLM.
+Specifically, after retrieving the top-$k$ documents using retrievals with a descending relevance score, we randomly shuffle the order of the retrieved documents and then input them to an LLM.
 In the second setting, we test how noisy documents impact the performance of LLMs.
@@
 However, when presented with noisy documents, all models experience a decline in performance.
-The performance drop for our self-reasoning framework is relatively minimal, which demonstrates that our proposed method is robust even when dealing with noisy documents.
+The performance drop in our self-reasoning framework is relatively minimal, demonstrating the robustness of our method even when handling noisy documents.
 Figure 4: Human citation quality evaluation vs. automatic citation evaluation on the long-form ASQA dataset.
-6.3 Citation analysis
+Citation Analysis
 As the automatic evaluation by the NLI model cannot detect partially supported citations, we discuss the analysis of citations with human evaluation in this section.
-Similarly to Liu et al.
-(2023), we conduct a human evaluation on two dimensions: 1) citation recall: annotators are given a statement and all documents that the statement refers to and are asked to judge whether the documents fully support the given statement; 2) citation precision: given a statement and one of its citations, annotators are asked to validate whether the citation fully supports, partially supports or does not support the statement.
-Each citation gets a precision score of 1 if the output sentence has a citation recall of 1 and this citation at least partially supports it.
-Details of human annotation can be found in Appendix §$\S$§A.8.
-As shown in Figure 4, we observe that the relative rankings by human evaluation align well with those from the automatic evaluation, and the human evaluation often yields a closely higher score when compared with the automatic evaluation.
-7 Conclusion
+Similarly to Liu, Zhang, and Liang (2023), we conduct a human evaluation on two dimensions: 1) citation recall: annotators are given a statement and all documents that the statement refers to and are asked to judge whether the documents fully support the given statement; 2) citation precision: given a statement and one of its citations, annotators are asked to validate whether the citation fully supports, partially supports or does not support the statement.
+As shown in Figure 4, the relative rankings by human evaluation align well with those from the automatic evaluation, and the human evaluation often yields a closely higher score when compared with the automatic evaluation.
+Details of human annotation can be found in the Appendix.
+Latency Analysis
+We also compared the inference latency of self-reasoning RAG with that of Self-RAG and GPT-4.
+The results show that our method maintains comparable latency to Self-RAG while delivering substantial performance gains.
+Detailed results are available in the Appendix.
+Conclusion
 RALMs can effectively enhance the performance of LLMs in handling knowledge-intensive tasks.
@@
 We conduct extensive experiments on four public datasets to demonstrate the superiority of our framework over existing state-of-the-art models.
-8 Limitations
-In this work, we mainly focus on improving the performance of RALMs with a self-reasoning framework on the task of open domain question answering and fact verification.
-Although we believe our framework can involve the distribution of real-world user questions, as we have evaluated in four public datasets, we acknowledge that we have not explored more challenging scenarios, such as multi-hop reasoning, code generation, and arithmetic reasoning.
-In future work, more challenging reasoning tasks, such as arithmetic reasoning, should be explored for the self-reasoning framework.
-We believe that our framework can effectively mitigate factual hallucinations in LLMs and improve the robustness of RALMs.
-However, there is still a risk that our method might generate hallucinations.
 Appendix A Appendix
@@
 Self-Reasoning (Ours) Y N N No Need 2K
-Self-RAG Asai et al. (2024) N N N No Need 145K (Generator)/ 46K (Critic)
-ReAct Yao et al. (2023) N N Y No Need No Need
-RECOMP Xu et al. (2023a) N Y Y 152K No Need
+Self-RAG (Asai et al. 2024) N N N No Need 145K (Generator)/ 46K (Critic)
+ReAct (Yao et al. 2023) N N Y No Need No Need
+RECOMP (Xu et al. 2024) N Y Y 152K No Need
 Table 4: Categorical comparisons with strong baseline models.
@@
 Train data for LLM indicates the number of training samples needed to train with LLMs.
-A.1 More Related Work of LMs for Reasoning
-One of the most well-known methods of using LLMs for reasoning is the Chain-of-Thought (CoT) Wei et al.
-(2022), which demonstrates the capability of LLMs to create their thinking process for problem-solving.
+More Related Work of LMs for Reasoning
+One of the most well-known methods of using LLMs for reasoning is the Chain-of-Thought (CoT) (Wei et al.
+2022), which demonstrates the capability of LLMs to create their thinking process for problem-solving.
 Zhou et al.
@@
 Recent works have extended beyond the internal reasoning ability of LLMs to include interactions with external tools (e.g., search engines or retrievers) for solving complex tasks.
-The ReAct Yao et al.
-(2023) presents an iterative paradigm to combine reasoning and acting with LLMs for tackling language reasoning and decision-making tasks.
+The ReAct (Yao et al.
+2023) presents an iterative paradigm to combine reasoning and acting with LLMs for tackling language reasoning and decision-making tasks.
 Xu et al.
-(2023b) introduces a framework to enable information retrieval and LLMs to interact with each other effectively with chain-of-query decomposition.
+(2023) introduces a framework to enable information retrieval and LLMs to interact with each other effectively with chain-of-query decomposition.
 Pan et al.
@@
 Different from the above works, which are mostly based on relatively large LLMs (e.g., ChatGPT), our proposed method focuses on enhancing smaller LLMs (e.g., LLaMA2) using only a limited number of samples to achieve high robustness and interpretability through single-step interaction.
-A.2 Instructions
+Instructions
 The instructions for GPT-4 to generate self-reasoning trajectories are shown in Figure 5 (the short-form and long-form QA tasks) and Figure 6 (the fact verification task).
 The words in the orange font are key fields that need to be generated.
-A.3 Datasets Description
+Datasets Description
 We conducted an extensive experimental evaluation of two short-form QA datasets, one long-form QA dataset, and a fact verification dataset.
-NaturalQuestion (NQ) Kwiatkowski et al.
-(2019) contains real user questions issued to the Google search and answers found from Wikipedia by the annotators.
+NaturalQuestion (NQ) (Kwiatkowski et al.
+2019) contains real user questions issued to the Google search and answers found from Wikipedia by the annotators.
 NQ is created to train and evaluate automated question answering systems.
-PopQA Mallen et al.
-(2023) is a large-scale open-domain question answering dataset, consisting of entity-centric QA pairs.
+PopQA (Mallen et al.
+2023) is a large-scale open-domain question answering dataset, consisting of entity-centric QA pairs.
 Each question is made by converting a knowledge triplet retrieved from Wikidata using a template.
 In this work, we use PopQA to evaluate performance in long-tail settings.
-ASQA Stelmakh et al.
-(2022) is a long-form factoid dataset, and most questions can be answered by Wikipedia.
-Each question originates from AmbigQA Min et al.
-(2020) and represents an ambiguous query that requires multiple short answers to cover various aspects.
+ASQA (Stelmakh et al.
+2022) is a long-form factoid dataset, and most questions can be answered by Wikipedia.
+Each question originates from AmbigQA (Min et al.
+2020) and represents an ambiguous query that requires multiple short answers to cover various aspects.
 The dataset provides a long-form answer that contains all short answers.
-FEVER Thorne et al.
-(2018) is a fact verification dataset that contains claims generated by rewriting sentences extracted from Wikipedia and subsequently verified without knowledge of the sentence from which they were derived.
+FEVER (Thorne et al.
+2018) is a fact verification dataset that contains claims generated by rewriting sentences extracted from Wikipedia and subsequently verified without knowledge of the sentence from which they were derived.
 The claims are classified as Supported, Refuted, or NotEnoughInfo.
-A.4 Experiment settings
+Experiment settings
 Training settings.
-During gradual learning, we fine-tune the LLaMA-2 Touvron et al.
-(2023) model with our self-reasoning framework for 3 epochs with a batch size set to 32, leveraging the DeepSpeed library Rasley et al.
-(2020) and the ZeRO optimizer, and we use parameter partitioning ZeRO stage 3 with float16 precision.
+During gradual learning, we fine-tune the LLaMA-2 (Touvron et al.
+2023) model with our self-reasoning framework for 3 epochs with a batch size set to 32, leveraging the DeepSpeed library (Rasley et al.
+2020) and the ZeRO optimizer, and we use parameter partitioning ZeRO stage 3 with float16 precision.
 The learning rate $r_{a}$ for the first stage is set to 5e-5, the learning rate $r_{b}$ for the second stage is set to 3e-5, and the learning rate $r_{c}$ for the final stage is set to 1e-5.
@@
 Inference settings.
-We use the vLLM  [^2] Codes are available at https://github.com/vllm-project/vllm framework Kwon et al.
-(2023) to accelerate the inference speed during inference.
+We use the vLLM  [^2] Codes are available at https://github.com/vllm-project/vllm framework (Kwon et al.
+2023) to accelerate the inference speed during inference.
 The codes follow the Apache-2.0 license agreement.
@@
 For the data generation quality control setting, the threshold for citation recall is set to 0.8, and the threshold of citation precision is set to 0.8.
-A.5 Categorical Comparisons
-We differentiate our method from existing strong baseline models by categorizing and comparing it across six dimensions, as presented in Table 3.
+Categorical Comparisons
+We differentiate our method from existing strong baseline models by categorizing and comparing it across five dimensions, as presented in Table 4.
 As illustrated in the table, our method can stand out in several key aspects.
@@
 This efficiency in training drastically lowers the resources and time needed, making our method both cost-effective and scalable for practical applications.
-A.6 Case Study
+Algorithm 1 Data Quality Control
+1:the origin self-reasoning dataset $\mathcal{D}_{0}$ generated by GPT-4
+2:filtered high quality self-reasoning dataset $\mathcal{D}_{sr}$
+3:Initialize the evaluation metrics program $P$
+4:Initialize the citation score tools $T$ (Gao et al. 2023b)
+5:for $i=1$ to $N$ do
+6:     Evaluate whether the answer of sample $d_{i}$ is correct using program $P$
+7:     if True and $d_{i}\in$ Long-form QA dataset  then
+8:         Compute citation recall score $s_{r}$ for sample $d_{i}$
+9:         Compute citation precision score $s_{p}$ for sample $d_{i}$
+10:         if  $s_{r}\geq\delta_{r}$ and $s_{p}\geq\delta_{p}$  then
+11:              Add the training sample to $\mathcal{D}_{l}$
+12:         end if
+13:     end if
+14:     if True and $d_{i}\notin$ Long-form QA dataset  then
+15:         Add the training sample to $\mathcal{D}_{s}$
+16:     end if
+17:end for
+18:$\mathcal{D}_{sr}=\mathcal{D}_{l}\cup\mathcal{D}_{s}$
+19:return $\mathcal{D}_{sr}$
+Case Study
 In our case study, as illustrated in Figure 7, we compare the responses generated by the raw LLM, the standard RALM (e.g.
@@
 As the case illustrated above, by leveraging relevant documents and focusing on contextual evidence, our self-reasoning framework can achieve a precise and well-supported answer, highlighting its utility and robustness in complex information retrieval tasks.
-A.7 More Analysis on Ablation Study
-Effectiveness of Gradual Learning
+More Analysis on Ablation Study
+Effectiveness of Gradual Learning.
 Further, we validate the effect of gradual learning.
@@
 As shown in Table 2, the performance decline can be observed in three datasets, suggesting that gradual learning can help improve performance.
-Effectiveness of Quality Control
+Effectiveness of Quality Control.
 The effect of quality control on data generation is also evaluated in our work.
@@
 As shown in Table 2, the substitution of unfiltered training data leads to the degradation of the model results.
-A.8 Human Evaluation
+Human Evaluation
 We randomly sample 100 examples from the ASQA dataset and annotate the outputs of selected models.
@@
 Each annotator is required to verify the citation recall and citation precision acorrding to our provided scheme.
-The annotation scheme is inspired by Gao et al.
-(2023b) as follows:
+The annotation scheme is inspired by (Gao et al.
+2023b) as follows:
 Citation Recall.
@@
 We calculate the overall recall score for the model by averaging the recall scores of all statements.
-Citation Precision
+Citation Precision.
 The annotators are shown the question $q$ and a statement $s_{i}$ and one of its citation ${c^{(k)}_{i}}\in C_{i}$.
@@
 Finally, we calculate the overall precision score for the model by averaging the precision scores of all statements.
+The Pseudo-code of Data Quality Control
+The details and pseudo-code of data quality control are illustrated in Algorithm 1.
+In the table, $\mathcal{D}_{0}$ is the origin self-reasoning dataset generated by GPT-4.
+$P$ is the program to evaluate the answer according to the metrics.
+$T$ is the off-the-shelf tool (Gao et al.
+2023b) to calculate the citation scores.
+$N$ is the number of total samples from the origin dataset, and $d_{i}$ is the $i$-th training sample.
+$s_{r}$ and $s_{c}$ are citation recall and precision scores for each sample respectively.
+$\delta_{r}$ and $\delta_{p}$ are thresholds of citation recall and precision respectively.
+$\mathcal{D}_{l}$ is filtered trajectory data for long-form QA, and $\mathcal{D}_{s}$ is filtered dataset for short-form QA and fact verification dataset.
+$\mathcal{D}_{sr}$ is the final high quality self-reasoning training dataset.
 Figure 5: The instructions for the GPT-4 to generate the self-reasoning trajectories for short-form and long-form QA tasks.
@@
 Figure 7: A Case Study. We present the self-reasoning trajectories generated by our framework during inference and demonstrate how they can logically generate the correct answer.
+Latency Analysis
+We conduct extra experiments to measure inference latency across our approach, Self-RAG, and GPT-4.
+The results demonstrated that our method achieved better performance while maintaining latency comparable to Self-RAG.
+The results of average inference latency per question for the NQ and ASQA datasets are shown in Table 5.
+Models NQ ASQA
+LLaMA2 0.19 1.92
+Self-RAG 2.08 2.38
+Self-Reasoning 2.19 2.32
+[…]
+Table 5: The latency experiment results on NQ and ASQA datasets. The average inference latency per question is evaluated using 7B-parameter models.